OpenStudy (anonymous):

6z +4 - 2z is greater than or equal to 3z - (5 + z)

OpenStudy (anonymous):

This is almost identical to the last problem I showed you. If you understand the last problem, you should be able to do this problem. If you cant do this problem, then redo the last problem.

OpenStudy (anonymous):

wouldn't you distribute first

jigglypuff314 (jigglypuff314):

yes, you would distribute first, then reduce, then simplify

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

do you subtract 2z from 6z

jigglypuff314 (jigglypuff314):

yes

jigglypuff314 (jigglypuff314):

\[6z+4-2z \ge 3z-(5+z) \rightarrow 6z+4-2z \ge 3z-5-z\] then \[6z+4-2z \ge 3z-5-z \rightarrow 4z + 4 \ge 2z - 5\]

OpenStudy (anonymous):

shouldn't you end up with 4z +4 is greater than or equal to 1z - (5 + z )

jigglypuff314 (jigglypuff314):

no?

OpenStudy (anonymous):

i did something wrong

jigglypuff314 (jigglypuff314):

6z+4−2z≥3z−(5+z) → 6z+4−2z≥3z−5−z 6z+4−2z≥3z−5−z → 4z+4≥2z−5

OpenStudy (texaschic101):

6z + 4 - 2z >= 3z - (5 +z) -- distribute through the parenthesis 6z + 4 - 2z >= 3z - 5 - z -- combine like terms 4z + 4 >= 2z - 5 -- subtract 2z from both sides 4z - 2z + 4 >= -5 -- subtract 4 from both sides 2z >= -5 -4 2z >= -9 z >= -9/2

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

anytime :)